# Exploring Multiple Strategies to Improve Multilingual Coreference Resolution in CorefUD

### v2

@@ Abstract
Coreference resolution, the task of identifying expressions in text that refer to the same entity, is a critical component in various natural language processing (NLP) applications. This paper presents our end-to-end neural coreference resolution system, utilizing the CorefUD 1.1 dataset, which spans 17 datasets across 12 languages. Our model is based on the end-to-end neural coreference resolution system by [1]. We first establish strong baseline models, including monolingual and cross-lingual variations, and then propose several extensions to enhance performance across diverse linguistic contexts. These extensions include cross-lingual training, incorporation of syntactic information, a Span2Head model for optimized headword prediction, and advanced singleton modeling. We also experiment with headword span representation and long-document modeling through overlapping segments. The proposed extensions, particularly the heads-only approach, singleton modeling, and long document prediction, significantly improve performance across most datasets. We also perform zero-shot cross-lingual experiments, highlighting the potential and limitations of cross-lingual transfer in coreference resolution. Our findings contribute to the development of robust and scalable coreference systems for multilingual coreference resolution. Finally, we evaluate our model on the CorefUD 1.1 test set and surpass the best model from the CRAC 2023 shared task of comparable size by a large margin. Our model is available on GitHub: https://github.com/ondfa/coref-multiling
@@ 1 Introduction
Coreference resolution is the task of identifying language expressions that refer to the same real-world entity (antecedent) within a text. These coreferential expressions can sometimes appear within a single sentence, but often, they are spread across multiple sentences. In some challenging cases, it is necessary to consider the entire document to determine whether two expressions refer to the same entity. The task can be divided into two main subtasks: identifying entity mentions and grouping these mentions based on the real-world entities they refer to. Coreference resolution is closely related to anaphora resolution, as discussed in [2]
@@
This paper describes our approach to multilingual coreference resolution. The task is based on the CorefUD dataset [3]. The CorefUD 1.1 corpus comprises 17 different datasets spanning 12 languages, all within a harmonized annotation scheme. As CorefUD is intended to extend Universal Dependencies with coreference annotations, all datasets within CorefUD are treebanks. For some languages, dependency annotations were provided by human annotators, while for others, these annotations were generated automatically using a parser. Coreference annotations are built upon these dependencies, meaning that mentions are represented as subtrees in the dependency tree and can be captured by their heads. In some datasets, there are non-treelet mentions—mentions that do not form a single subtree—but even for these, a single headword is selected. Notable differences exist between the datasets, with one of the most prominent being the presence of singletons. Singletons are clusters that contain only one mention and do not participate in any coreference relation, yet they are still annotated as mentions. For further details on the dataset, refer to [3] or [5]. We use the primary evaluation metric from CRAC 2023 Multilingual Coreference Resolution Shared Task. The metric ignores singletons. Building upon the baseline model proposed by [1], we introduce several novel extensions aimed at enhancing the performance of coreference resolution across multiple languages and datasets within the CorefUD collection. Our primary goal is to develop a universal model capable of handling the diverse and complex nature of these datasets effectively.
@@
In response to the evaluation metrics used in the CRAC 2022 Shared Task, we also propose the Span2head model, which simplifies span-based mention representation to headword prediction. This adjustment aligns the model’s output more closely with the evaluation criteria, where we consider two mentions the same if they share the same head.
Additionally, we explore the potential of head representations as a simplified approach to coreference resolution. By modeling mentions solely based on their syntactic heads, we reduce the computational complexity from quadratic to linear, making the model more efficient and less prone to errors, particularly in the case of long and complex mentions.
@@
Through these extensions, our model aims to advance the state of multilingual coreference resolution by addressing the specific challenges posed by the diverse datasets in the CorefUD collection, ultimately contributing to more accurate and generalizable coreference systems.
@@ 2.1 End-to-end Neural Coreference Resolution Models
The model estimates the probability of a mention corefering with the antecedent in a document . Since the model adopts the end-to-end approach, the mentions are identified together with the coreference links. We consider every continuous sequence of words as a mention . Therefore, we work with possible mentions, where is the number of words in a document .
We model the score of a mention corefering with an antecedent as a combination of two types of scores and . The is a score of a sequence of words (spans) being a mention. The score is the score of a span being an antecedent of span . The scores are combined as a sum of , and as follows:
where is an empty (also referred as dummy) antecedent. If a mention candidate is aligned to the dummy antecedent and it is not an antecedent of any other mention (so the mention should be in the cluster with dummy antecedent only), it is not considered a valid mention. As a result, every valid mention has to be in a cluster with at least one other mention, so the model cannot predict singletons. Both scores and are estimated with a feed-forward neural network over the BERT-based encoder.
The probability of an antecedent can be expressed as the normalization over all possible antecedents for a mention :
The formula for all antecedents uses a product of multinomials of all individual antecedents:
In the training phase, we maximize the marginal log-likelihood of all correct antecedents:
where GOLD() is the set of spans (mentions) in the training data that are antecedents of mention .
@@
This end-to-end model has proven to be highly successful, especially on datasets where singletons are not annotated (e.g., CoNLL 2012/OntoNotes), a scenario where traditional two-stage models tend to struggle. The model was later extended by [7] to the higher-order coreference resolution (CR) through iterative refining of span representations with a gated attention mechanism. They further optimize pro model speed by pruning the mentions with simple scorer at the first step and rescoring only top spans with the precise scorer and the second step. [8] uses the model from [7] but with BERT as an encoder.
@@ 2.2 Word-level Coreference Resolution
[10] proposed a word-level coreference resolution approach based on the model by [6]. Their method reduces the computational complexity by reducing the mention space. Instead of iterating over all possible spans they first map each gold mention to a single word—specifically, the headword, which they select based on the syntax tree (as the node closest to the root of the tree). Antecedent prediction is then performed on the word level. In the next step, they use a span extraction model to predict the original spans from headwords. This model is trained to classify each word in the sentence together with a headword to decide, whether the word is the start token or end token of the span corresponding to the headword. During training the span extraction model is trained simultaneously with with the antecedent prediction as another classification head.
Later [11] proposed a modification in headword selection, suggesting that when a coordinate conjunction appears within an entity, it should be made the head to resolve span ambiguity. For example, in the sentence "Tom and Mary are playing," the entities "Tom" and "Tom, and Mary" would typically share the same head in the original approach. Their modification eliminates this ambiguity.
@@ 2.3 CRAC Shared Task on Multilingual Coreference Resolution
CRAC Shared Task on Multilingual Coreference Resolution (CRAC-coref) [12] is an annual shared task that began in 2022 and is built upon the CorefUD collection. In fact, this paper is an extension of CRAC22-coref [13] participant system [14], which is based on [15]
Best-performing system in CRAC 2022 was submitted by [16]. It is a two-stage model with a joined BERT-like encoder. In the mention detection stage, they use extended BIO111BIO stands for Beginning, Inside, and Outside. It is a tagging format to label each token of a text as being the beginning (B), inside (I), or outside (O) of an entity. tagging via stack manipulation instructions to solve overlapping mentions. Coreference resolution is then done by antecedent prediction head, a component responsible for identifying the most likely antecedent of each mention, which is very similar to the one used in end-to-end neural CR [6]. They use a concatenation of start and end tokens as mention representation and classify mention pairs to find the best antecedent of each mention. The schema of the CorPipe model is shown in Figure 2.
@@ 3 Datasets & Metrics
For our experiments, we use the CorefUD 1.1 dataset. The collection consists of 17 datasets in 12 different languages. Table 1 provides a summary of these datasets and their respective sizes.
3.1 Dataset Differences
@@
As highlighted by [3] significant differences remain among the individual datasets within CorefUD, not only in data distribution but also in annotation practices. In this section, we discuss these differences that are particularly relevant to our model and experiments.
3.1.1 Empty Nodes
@@
Some datasets contain empty nodes. These are virtual nodes added to a syntactic tree during the annotation process (and also into the text). These nodes refer to unexpressed entities; entities that are not explicitly mentioned in the text but are referenced. Adding empty nodes can make it hard for a general-purpose pre-trained model to to accurately interpret the meaning of the text. We cannot discard them from the datasets because most of them are part of a coreference cluster. Entity mentions that span only the empty nodes are called zero mentions.
3.1.2 Singletons
@@
Another key difference between the datasets lies in the annotation of singletons. Singletons are entities with only a single mention throughout the document so these mentions have no coreference relations. In some datasets, singletons are annotated, while in others they are not. That means that when there is an entity with only a single mention, it is not considered an entity during annotation. Table 2 provides the percentage of singleton entities across the datasets (entities with a length of 1).
3.2 Dataset Lengths
Tables 1, 2, and 3 show basic statistics of all the datasets in CorefUD collection. All these tables are taken from the CorefUD paper [5]. Table 1 shows the basic sizes of the datasets and portion of the train/dev/test split. Column empty shows the number of empty nodes. Table 2 shows statistics about entities. The length of an entity means the number of its mentions thus entities with length 1 are singletons. Table 3 shows statistics of mentions. Here the length of a mention means the number of its words. The last three columns show the percentage of specific mentions. w/empty are mentions with at least one empty node, w/gap are mentions with at least one gap e.g. discontinuous mentions. non-tree are mentions which do not form a single subtree of the syntax tree.
total size | division [%] | ||||||
CorefUD dataset | docs | sents | words | empty | train | dev | test |
Catalan-AnCora | 1298 | 13,613 | 429,313 | 6,377 | 77.5 | 11.4 | 11.1 |
Czech-PCEDT | 2312 | 49,208 | 1,155,755 | 35,844 | 80.9 | 14.2 | 4.9 |
Czech-PDT | 3165 | 49,428 | 834,720 | 22,389 | 78.3 | 10.6 | 11.1 |
English-GUM | 195 | 10,761 | 187,416 | 99 | 78.9 | 10.5 | 10.6 |
English-ParCorFull | 19 | 543 | 10,798 | 0 | 81.2 | 10.7 | 8.1 |
French-Democrat | 126 | 13,057 | 284,883 | 0 | 80.1 | 9.9 | 10.0 |
German-ParCorFull | 19 | 543 | 10,602 | 0 | 81.6 | 10.4 | 8.1 |
German-PotsdamCC | 176 | 2,238 | 33,222 | 0 | 80.3 | 10.2 | 9.5 |
Hungarian-KorKor | 94 | 1,351 | 24,568 | 1,988 | 79.2 | 10.3 | 10.5 |
Hungarian-SzegedKoref | 400 | 8,820 | 123,968 | 4,857 | 81.1 | 9.6 | 9.3 |
Lithuanian-LCC | 100 | 1,714 | 37,014 | 0 | 81.3 | 9.1 | 9.6 |
Norwegian-BokmaalNARC | 346 | 15,742 | 245,515 | 0 | 82.8 | 8.8 | 8.4 |
Norwegian-NynorskNARC | 394 | 12,481 | 206,660 | 0 | 83.6 | 8.7 | 7.7 |
Polish-PCC | 1828 | 35,874 | 538,885 | 470 | 80.1 | 10.0 | 9.9 |
Russian-RuCor | 181 | 9,035 | 156,636 | 0 | 78.9 | 13.5 | 7.6 |
Spanish-AnCora | 1356 | 14,159 | 458,418 | 8,112 | 80.0 | 10.0 | 10.0 |
Turkish-ITCC | 24 | 4,733 | 55,341 | 0 | 81.5 | 8.8 | 9.7 |
Entities | distribution of lengths | ||||||||
CorefUD dataset | total | per 1k | length | 1 | 2 | 3 | 4 | 5+ | |
count | words | max | avg. | [%] | [%] | [%] | [%] | [%] | |
Catalan-AnCora | 18,030 | 42 | 101 | 3.5 | 2.6 | 54.0 | 18.1 | 8.6 | 16.8 |
Czech-PCEDT | 52,721 | 46 | 236 | 3.3 | 6.6 | 59.6 | 14.7 | 6.3 | 12.7 |
Czech-PDT | 78,747 | 94 | 175 | 2.4 | 40.8 | 35.1 | 10.4 | 4.9 | 8.9 |
English-GUM | 27,757 | 148 | 131 | 1.9 | 73.8 | 14.2 | 4.9 | 2.2 | 4.9 |
English-ParCorFull | 202 | 19 | 38 | 4.2 | 6.9 | 54.0 | 13.9 | 5.9 | 19.3 |
French-Democrat | 39,023 | 137 | 895 | 2.0 | 81.6 | 10.7 | 3.0 | 1.4 | 3.3 |
German-ParCorFull | 259 | 24 | 43 | 3.5 | 6.2 | 64.9 | 11.6 | 5.0 | 12.4 |
German-PotsdamCC | 3,752 | 113 | 15 | 1.4 | 76.5 | 13.9 | 5.0 | 1.8 | 2.7 |
Hungarian-KorKor | 1,134 | 46 | 41 | 3.6 | 0.9 | 55.1 | 17.1 | 9.0 | 17.9 |
Hungarian-SzegedKoref | 5,182 | 42 | 36 | 3.0 | 8.0 | 51.1 | 19.0 | 9.1 | 12.9 |
Lithuanian-LCC | 1,224 | 33 | 23 | 3.7 | 11.2 | 45.3 | 11.8 | 8.2 | 23.5 |
Norwegian-BokmaalNARC | 53,357 | 217 | 298 | 1.4 | 89.4 | 5.4 | 1.9 | 1.0 | 2.4 |
Norwegian-NynorskNARC | 44,847 | 217 | 84 | 1.4 | 88.7 | 5.5 | 2.1 | 1.1 | 2.7 |
Polish-PCC | 127,688 | 237 | 135 | 1.5 | 82.6 | 9.8 | 2.9 | 1.4 | 3.2 |
Russian-RuCor | 3,636 | 23 | 141 | 4.5 | 3.3 | 53.7 | 15.6 | 6.9 | 20.5 |
Spanish-AnCora | 20,115 | 44 | 110 | 3.5 | 3.3 | 53.7 | 17.0 | 8.8 | 17.1 |
Turkish-ITCC | 690 | 12 | 66 | 5.3 | 0.0 | 39.0 | 19.6 | 10.7 | 30.7 |

### v3

@@ Abstract
Coreference resolution, the task of identifying expressions in text that refer to the same entity, is a critical component in various natural language processing applications. This paper presents a novel end-to-end neural coreference resolution system, utilizing the CorefUD 1.1 dataset, which spans 17 datasets across 12 languages. The proposed model is based on the standard end-to-end neural coreference resolution system. We first establish baseline models, including monolingual and cross-lingual variations, and then propose several extensions to enhance performance across diverse linguistic contexts. These extensions include cross-lingual training, incorporation of syntactic information, a Span2Head model for optimized headword prediction, and advanced singleton modeling. We also experiment with headword span representation and long-documents modeling through overlapping segments. The proposed extensions, particularly the heads-only approach, singleton modeling, and long document prediction, significantly improve performance across most datasets. We also perform zero-shot cross-lingual experiments, highlighting the potential and limitations of cross-lingual transfer in coreference resolution. Our findings contribute to the development of robust and scalable coreference systems for multilingual coreference resolution. Finally, we evaluate our model on the CorefUD 1.1 test set and surpass the best model from the CRAC 2023 shared task of comparable size by a large margin.
@@ 1 Introduction
Coreference resolution is the task of identifying which expressions in a text refer to the same entity in the real world. The task can be divided into two subtasks. Identify entity mentions and group them according to the entities they refer to. Anaphora resolution is closely related to coreference resolution. Anaphora is a contextual reference to an expression in the previous context. Most of the anaphoric relations are also coreferential, but it does not necessarily be so. For example, in the sentence:
This apple is mine. Please, take another one,
the two mentions are anaphoric, but they are not coreferential (they refer to different apples). Some datasets contain only anaphora, some only coreferences, and others contain annotations of both. Anaphoric relations are based more on the syntax of a sentence, whereas coreference relations are more about semantics. Coreferences are usually annotated in terms of coreference clusters or in terms of coreference chains. Coreference clusters group together all the mentions of the same entity, no matter their order. In the coreference chain, each mention is aligned with its direct antecedent (closest previous mention of the same entity). Although significant advancements have been made in coreference resolution for English, developing robust and effective models for multiple languages remains a challenging task due to linguistic diversity and the availability of resources.
@@
This paper proposes a novel approach to multilingual coreference resolution. We use the CorefUD 1.1 dataset [1]. The CorefUD 1.1 is a harmonized corpus with 17 different datasets spanning 12 languages. The dataset is intended to extend Universal Dependencies [2] with coreference annotations. Therefore, all datasets within CorefUD are treebanks. Coreference annotations are built upon these dependencies, meaning that mentions are represented as subtrees in the dependency tree and can be captured by their heads. Moreover, the evaluation metric used with this dataset on CRAC 2023 Multilingual Coreference Resolution Shared Task uses head-match evaluation. This means that the predicted mention is considered the same as the gold mention if they share the same head. Notable differences exist between the datasets, with one of the most prominent being the presence of singletons. Singletons are clusters that contain only one mention and do not participate in any coreference relation, yet they are still annotated as mentions. For further details on the dataset, refer to [1].
Building upon the baseline model proposed by [3], we introduce several novel extensions aimed at enhancing the performance of coreference resolution across multiple languages and datasets within the CorefUD collection. Our primary goal is to develop a universal model capable of handling the diverse and complex nature of these datasets effectively.
@@
In response to the evaluation metrics used in the CRAC 2022 Shared Task, we also evaluate the Span2head model, which simplifies span-based mention representation to headword prediction. This adjustment aligns the model’s output more closely with the evaluation criteria, where we consider two mentions the same if they share the same head.
Additionally, we explore the potential of head representations as a simplified approach to coreference resolution. Similarly to word-level coreference [4], by modeling mentions solely based on their syntactic heads, we reduce the computational complexity from quadratic to linear, making the model more efficient and less prone to errors, particularly in the case of long and complex mentions.
@@
We experimentally evaluate all proposed extensions to find the best configuration for a joined model for all the datasets in CorefUD. Through these extensions, our model aims to advance the state of multilingual coreference resolution by addressing the specific challenges posed by the diverse datasets in the CorefUD collection, ultimately contributing to more accurate and generalizable coreference systems.
The main contributions of the paper are the following:
We propose several extensions of a standard model that lead to significant improvements. The most prominent extensions include singletons modeling, headword mention representation, and segments overlapping for long coreference prediction.
We provide a detailed evaluation of all proposed extensions.
We perform the analysis of knowledge transferability between the datasets with joined training and cross-lingual zero-shot experiments.
We demonstrated that the proposed model significantly surpasses the state of the art with a comparable model size.
We trained a single model that has a great performance on all the datasets in the CorefUD collection.
Our source codes are publicly available on GitHub for subsequent research: https://github.com/ondfa/coref-multiling.
@@ 2.1 End-to-end Neural Coreference Resolution Models
The model estimates the probability of a mention corefering with the antecedent in a document . The model is trained in an end-to-end manner, so every continuous span up to the maximum length is considered as a mention. Therefore, we work with possible mentions, where is the number of words in a document .
is estimated from two scores: The first is the mention score of both mention and antecedent , which express the likelihood of a span being a valid mention. The second is the binary antecedent score of being an antecedent of mention .
All three scores are summed together as follows:
where is an empty (also referred as dummy) antecedent. If a mention candidate is aligned to the dummy antecedent and it is not an antecedent of any other mention (so the mention should be in the cluster with dummy antecedent only), it is not considered a valid mention. As a result, every valid mention has to be in a cluster with at least one other mention, so the model cannot predict singletons.
The scores are then normalized with the softmax function to estimate the probability:
The scores are estimated by FFNN with one output neural without output activation as follows:
where is a vector representation of span obtained from text encoder (originally LSTM).
where represents pairwise features lake distance between spans.
Vector representation of a span is optained from a concatenation of various features as follows:
where is the whole input and , are start and end token respectively. is the sum of all the span tokens weighted with the attention mechanism [6]. is an additional feature; span length.
In the training phase, we iterate all mentions predicted by a model and maximize the marginal probability of all its correct antecedents as shown in the following equation:
where GOLD() is the set of all gold antecedents of mention (in the training data annotation).
Theoretically, we should go over all possible mentions, but in practice, the mentions are prefiltered based on a unary mention score to reduce computational complexity.
@@
This end-to-end model has proven to be highly successful, especially on datasets where singletons are not annotated (e.g., CoNLL 2012/OntoNotes), a scenario where traditional two-stage models tend to struggle. The model was later extended by [7] to the higher-order coreference resolution (CR) through iterative refining of span representations with gated attention mechanism. They further optimize pro model speed by pruning the mentions with simple scorer at the first step and rescoring only top spans with the precise scorer in the second step. Formally, they add the coarse coreference score: , which is used for antecedent pruning before computing the fine coreference score . The method is called coarse-to-fine (c2f) coreference resolution. Authors of [8] use the model from [7] but with BERT as an encoder.
@@ 2.2 Word-level Coreference Resolution
Authors of [4] proposed a word-level coreference resolution approach based on the model by [5]. Their method reduces the computational complexity by reducing the mention space. Instead of iterating over all possible spans, they first map each gold mention to a single word—specifically, the headword, which they select based on the syntax tree (as the node closest to the root of the tree). Antecedent prediction is then performed on the word level. In the next step, they use a span extraction model to predict the original spans from headwords. This model is trained to classify each word in the sentence together with a headword to decide whether the word is the start token or end token of the span corresponding to the headword. During training, the span extraction model is trained simultaneously with the antecedent prediction as another classification head. The rest of the model is standard coarse-to-fine coreference resolution.
Later in [10], the authors proposed a modification in headword selection, suggesting that when a coordinate conjunction appears within an entity, it should be made the head to resolve span ambiguity. For example, in the sentence "Tom and Mary are playing," the entities "Tom" and "Tom, and Mary" would typically share the same head in the original approach. Their modification eliminates this ambiguity.
@@ 2.3 CRAC Shared Task on Multilingual Coreference Resolution
CRAC Shared Task on Multilingual Coreference Resolution (CRAC-coref) [11] is an annual shared task that began in 2022 and is built upon the CorefUD collection. In fact, this paper is an extension of CRAC22-coref [12] participant system [13], which is based on [14].
The best-performing system in CRAC 2022 was submitted by [15]. It is a two-stage model with a joined BERT-like encoder. In the mention detection stage, they use extended BIO111BIO stands for Beginning, Inside, and Outside. It is a tagging format to label each token of a text as being the beginning (B), inside (I), or outside (O) of an entity. tagging via stack manipulation instructions to solve overlapping mentions. Coreference resolution is then done by antecedent prediction head, a component responsible for identifying the most likely antecedent of each mention, which is very similar to the one used in end-to-end neural CR [5]. They use a concatenation of start and end tokens as mention representation and classify mention pairs to find the best antecedent of each mention. The schema of the CorPipe model is shown in Figure 2.
@@ 3 Datasets & Metrics
For the evaluation of the proposed model, we use the CorefUD 1.1 dataset. The collection consists of 17 datasets in 12 different languages. All datasets in the collection are converted into a harmonized form.
The primary goal of CorefUD is to have as many datasets as possible in the same annotation scheme and enable cross-lingual training. Several important design choices have been made by the authors of CorefUD:
As the collection is the extension of Universal Dependencies [2], all the datasets in the CorefUD collection are theebanks. Some of them were annotated in a treebank form (e.g., Czech-PDT), and the others were parsed during the harmonization procedure.
Coreference entities are represented as clusters of mentions.
@@
As highlighted in [1], significant differences remain among individual datasets within CorefUD, not only in data distribution but also in annotation practices. In this section, we discuss the differences that are particularly relevant to our model and experiments.
Zero mentions and empty nodes:
@@
Some datasets contain zero mentions nodes. These are mentions that are not explicitly mentioned in the text (cannot be mapped to a word or phrase), but they are implicitly referred to. In CorefUD, zero mantions are captured in empty nodes; virtual nodes are added to a syntactic tree during the annotation process, and they are not mapped to any word. Adding empty nodes as some special words can make it hard for a general-purpose pre-trained model to interpret the meaning of the text accurately. We cannot discard them from the datasets because most of them are part of a coreference cluster. The number of empty nodes in each dataset is shown in Table 1, and the number of zero mentions can be found in Table 3.
Singletons:
@@
Another critical difference between the datasets lies in the annotation of singletons. Singletons are entities with only a single mention throughout the document, so these mentions have no coreference relations. In some datasets, singletons are annotated, while in others, they are not. That means that when there is an entity with only a single mention, it is not considered an entity during annotation. Table 2 provides the percentage of singleton entities across the datasets (entities with a length of 1).
The CorefUD project is still evolving, and every year, there are some new datasets. The data is also moving towards a realistic scenario. In the very last version, 1.2, test data are practically just plain text (parsed with UDPipe, but no manual annotation is provided).
total size division [%] CorefUD dataset docs sents words empty nodes train dev test Catalan-AnCora 1298 13,613 429,313 6,377 77.5 11.4 11.1 Czech-PCEDT 2312 49,208 1,155,755 35,844 80.9 14.2 4.9 Czech-PDT 3165 49,428 834,720 22,389 78.3 10.6 11.1 English-GUM 195 10,761 187,416 99 78.9 10.5 10.6 English-ParCorFull 19 543 10,798 0 81.2 10.7 8.1 French-Democrat 126 13,057 284,883 0 80.1 9.9 10.0 German-ParCorFull 19 543 10,602 0 81.6 10.4 8.1 German-PotsdamCC 176 2,238 33,222 0 80.3 10.2 9.5 Hungarian-KorKor 94 1,351 24,568 1,988 79.2 10.3 10.5 Hungarian-SzegedKoref 400 8,820 123,968 4,857 81.1 9.6 9.3 Lithuanian-LCC 100 1,714 37,014 0 81.3 9.1 9.6 Norwegian-BokmaalNARC 346 15,742 245,515 0 82.8 8.8 8.4 Norwegian-NynorskNARC 394 12,481 206,660 0 83.6 8.7 7.7 Polish-PCC 1828 35,874 538,885 470 80.1 10.0 9.9 Russian-RuCor 181 9,035 156,636 0 78.9 13.5 7.6 Spanish-AnCora 1356 14,159 458,418 8,112 80.0 10.0 10.0 Turkish-ITCC 24 4,733 55,341 0 81.5 8.8 9.7
Entities distribution of lengths CorefUD dataset total per 1k length 1 2 3 4 5+ count words max avg. [%] [%] [%] [%] [%] Catalan-AnCora 18,030 42 101 3.5 2.6 54.0 18.1 8.6 16.8 Czech-PCEDT 52,721 46 236 3.3 6.6 59.6 14.7 6.3 12.7 Czech-PDT 78,747 94 175 2.4 40.8 35.1 10.4 4.9 8.9 English-GUM 27,757 148 131 1.9 73.8 14.2 4.9 2.2 4.9 English-ParCorFull 202 19 38 4.2 6.9 54.0 13.9 5.9 19.3 French-Democrat 39,023 137 895 2.0 81.6 10.7 3.0 1.4 3.3 German-ParCorFull 259 24 43 3.5 6.2 64.9 11.6 5.0 12.4 German-PotsdamCC 3,752 113 15 1.4 76.5 13.9 5.0 1.8 2.7 Hungarian-KorKor 1,134 46 41 3.6 0.9 55.1 17.1 9.0 17.9 Hungarian-SzegedKoref 5,182 42 36 3.0 8.0 51.1 19.0 9.1 12.9 Lithuanian-LCC 1,224 33 23 3.7 11.2 45.3 11.8 8.2 23.5 Norwegian-BokmaalNARC 53,357 217 298 1.4 89.4 5.4 1.9 1.0 2.4 Norwegian-NynorskNARC 44,847 217 84 1.4 88.7 5.5 2.1 1.1 2.7 Polish-PCC 127,688 237 135 1.5 82.6 9.8 2.9 1.4 3.2 Russian-RuCor 3,636 23 141 4.5 3.3 53.7 15.6 6.9 20.5 Spanish-AnCora 20,115 44 110 3.5 3.3 53.7 17.0 8.8 17.1 Turkish-ITCC 690 12 66 5.3 0.0 39.0 19.6 10.7 30.7